# Glimpse: Generalized Local Imaging with MLPs

### v2

@@ Abstract
Deep learning has become the state-of-the-art approach to medical tomographic imaging. A common approach is to feed the result of a simple inversion, for example the backprojection, to a (multiscale) convolutional neural network (CNN), which then computes the final reconstruction. Despite strong results on in-distribution test data similar to the training data, they overfit certain large-scale structures which leads to poor generalization on out-of-distribution (OOD) samples. Moreover, their memory complexity and training time scale unfavorably with image resolution, making them impractical for application at realistic clinical resolutions, especially in 3D. A standard U-Net requires a substantial 140GB of memory and 2600 seconds per epoch on a research-grade GPU when training on 1024 $\times$ 1024 images with batch size 64. In this paper, we introduce Glimpse, a local processing neural network for computed tomography which reconstructs a pixel value by processing only the measurements associated with the neighborhood of the pixel with a simple multi-layer perception (MLP). While achieving performance comparable to or better than successful CNNs like U-Net on in-distribution test data, Glimpse significantly outperforms them on OOD samples while maintaining a memory footprint almost independent of image resolution; 5GB memory suffices to train on 1024 $\times$ 1024 images, with each epoch requires 420 seconds. Because we built Glimpse to be fully differentiable it can also be used as a plug-in component of arbitrary deep learning architectures, enabling feats such as correction of miscalibrated projection orientations.
Index Terms:
Deep Learning, Computed Tomography, MLP, Uncalibrated Imaging
I Introduction
@@
Convolutional neural networks (CNNs) have become the standard approach for tomographic image reconstruction [1]. U-Net [2] has emerged as an architecture underpinning numerous deep learning reconstruction methods, applied with great success to a variety of imaging problems including computed tomography (CT) [3], magnetic resonance imaging (MRI) [4] and photoacoustic tomography [5]. Its success is often attributed to its particular multi-scale architecture [6].
At the same time, certain aspects of multi-scale CNNs complicate their application to real problems. Despite good performance on in-distribution test images similar to the training data, they often overfit specific image content resulting in poor generalization to distribution shifts in image content and sensing as shown in this paper. Model-based networks attempt to address this drawback by integrating the forward and adjoint operators into multiple network layers or iterations [7, 8, 9, 10, 11, 12]. However, the required memory for CNNs directly scales with image resolution [13]. For instance, the widely used U-Net requires a substantial 140GB memory and 2600 seconds per epoch when training on 1024 $\times$ 1024 images using two Tesla A100 GPUs. This latter drawback is further exacerbated with model-based networks such as learned primal-dual (LPD) [8], which achieves strong performance but requires over 80GB memory and very long training time even at a lower resolution of $512\times 512$. This increased memory demand is due to the repeated application of the forward model and its adjoint in forward and backward passes of the neural network. This makes standard CNN-based pipelines impractical for real-world scenarios involving resolutions beyond $512\times 512$.
@@
To better understand the mechanics behind the poor generalization of U-Net-like CNNs which compute the reconstruction from filtered backprojections (FBP) [14], we designed an experiment as follows. Figure 1 shows an object with a point source, its sparse view sinogram measurements with sinusoidal support, and the FBP reconstruction. It is evident that the FBP is supported over the entire field of view. This observation raises the question of the ideal receptive field size for CNNs like U-Net: a large receptive field may be beneficial to capture all information correlated with the value of a target pixel [15, 16].
However, models with large receptive fields often overfit specific image content in training data which leads to poor generalization on out-of-distribution samples [17]. Indeed, Figure 2 shows that while U-Net produces good results when tested on in-distribution data similar to training data (here chest images), it performs poorly on out-of-distribution (here brain images). This makes CNNs like U-Net problematic in domains such as medical imaging where robustness over distribution shifts and other uncertain and variable factors is of relevance [18].
@@
Figure 2: Performance of U-Net [2] trained on chest images: evaluation on in-distribution test data (chest samples) and OOD brain samples shows that the large receptive field of U-Net hinders its ability to generalize on OOD samples, with its PSNR even falling below that of FBP reconstruction.
In this paper, we introduce a new deep learning imaging architecture termed Glimpse—a simple local processing neural network adapted to the geometry of computed tomography. As shown in Figure 3, to recover the image intensity at a given target pixel, we use a MLP that takes only the local sinogram measurements associated with this pixel and its neighbors. There is no backprojection step. This localization results in robust performance, particularly when dealing with OOD data.
Figure 3: Glimpse; a single MLP processes the measurements associated with the pixel $(x,y)$ and its neighbors extracted from the sinogram. This local processing network has promising performance on OOD data while being computationally efficient all due to its locality.
At the same time, this design makes Glimpse highly computationally efficient; it permits training on mini-batches of both pixels and objects. This flexibility leads to fast and efficient training, requiring a small, fixed amount of memory almost independent from the image resolution. This allows training Glimpse on large, realistic images in resolution $1024\times 1024$ and beyond.
We built Glimpse to be fully differentiable, all the way down to the sensing and integration geometry. This has several advantages over the standard CNN-based architectures. For instance, most methods for CT image recovery strongly rely on the sensor geometry information encoded in the forward operator, whether explicitly, as seen in methods like FBP [14], SART [19], LGS [7], and LPD [8] or implicitly as used in U-Net [2] when taking FBP as input. This fixed geometry is a problem when faced with uncertainties in calibration or blind inversion problems where the sensor geometry information is entirely unavailable. While such uncertainties might degrade the quality of reconstructions of the standard methods [20, 21], our differentiable architecture allows the optimization of projection angles which can estimate the right projection angles and improve the quality of reconstructions.
II Related Work
@@
Model-based vs Model-free Inversion. There are two major classes of deep-learning-based approaches to CT: model-based and model-free inversion. In the model-based approach, neural networks process raw sinograms and map them to the desired CT images while the Radon transform is integrated into multiple network layers or iterations [7, 8, 11]. These methods perform remarkably well across various inverse problems, but they are computationally expensive, especially during training [13]. The high computational cost is due, among other factors, to the repeated application of the Radon transform and its adjoint in forward and backward passes of the neural networks.
@@
In contrast, model-free approaches offer a computationally cheaper alternative. The Radon transform (or its adjoint) is only used once in FBP computation before the neural network [3, 22, 23]. However, these models often require deep networks with a large receptive field to leverage the information delocalized across the FBP image. Recently, Hamoud et al. [16] used a measurement rearrangement technique to stratify backprojected features by angle and thus enable the use of smaller, shallower CNNs.
MLPs for Imaging A multi-layer perceptron (MLP) is a fundamental neural network architecture used in a great variety of applications. Recently, vision transformers [24] and MLP-mixers [25] have shown promising performance in various computer vision tasks like image classification [26] and image restoration [27]. While, unlike CNNs, a vanilla MLP lacks a good inductive bias for imaging, in particular translation equivariance, vision transformers and MLP-mixers restore it by processing patches instead of entire images [28]. However, these strategies require large datasets and networks to achieve performance comparable with CNNs. In our work, we propose a differentiable local processing network for CT imaging, demonstrating that even a small MLP can achieve performance on par with or even surpassing that of popular image-to-image CNNs.
@@
Uncalibrated CT Imaging. In CT imaging, the acquisition operator is often known but an insufficient number of measurements is obtained. This may occur when a reduced number of projections is used to minimize radiation exposure or shorten acquisition time (sparse view) or when only a limited cone of projection angles may be used (limited view). In certain situations, the acquisition operator is only partially or approximately known. Neglecting this uncertainty can result in a significant drop in the quality of the reconstructions [20]. To tackle this challenge, total least squares approaches have been developed, involving the perturbation of an assumed forward operator [29, 30, 31]. Recently, Gupta et al. [21] used autodifferentiation and gradient descent to estimate the uncalibrated forward operator in a self-consistent manner.
III Computed Tomography
@@
CT imaging [32] plays an important role in many applications including medical diagnosis [33], industrial testing [34], and security [35]. We consider 2D computed tomography where the image of interest $f({\mathbf{x}})$ with size $D\times D$ is reconstructed from measurements of (X-ray) attenuation. The forward model is the Radon transform $Rf$ which computes integrals of $f({\mathbf{x}})$ along lines $L$,
@@
${\mathbf{s}}={\mathbf{A}}{\mathbf{f}}+{\mathbf{n}}.$ (5)
@@
${\mathbf{f}}^{\text{FBP}}_{x,y}=\sum_{m=1}^{r}\tilde{{\mathbf{s}}}[y\cos(\alpha_{m})-x\sin(\alpha_{m}),m],$ (6)
where ${\mathbf{f}}^{\text{FBP}}\in\mathbb{R}^{N\times N}$ is the FBP reconstruction, $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$, ${\mathbf{h}}$ is a certain high-pass filter, * denotes the convolution and linear interpolation is used in (6) for evaluating $\tilde{{\mathbf{s}}}[x,\cdot]$ when $x$ is not an integer. As shown in Proposition 2 in Appendix -C, while the Ram–Lak filter is the optimal choice for ${\mathbf{h}}$ in the case of noise-free complete measurements, it can amplify the noise in real-world noisy measurements, leading to poor reconstruction.
With measurement noise and an incomplete collection of projections, tomographic image reconstruction from a sinogram becomes an ill-posed inverse problem that requires an image prior. In the following section, we introduce our proposed method, Glimpse, designed so that it respects the geometry of CT imaging.
III-A Glimpse: Local Imaging with MLPs
@@
$\text{SIN}_{x,y}[m]={\mathbf{s}}[y\cos(\alpha_{m})-x\sin(\alpha_{m}),m].$ (7)
Similar to (6), we can use interpolation to evaluate ${\mathbf{s}}[x,\cdot]$ for non-integer $x$. This localization is formally captured by the following proposition.
@@
The proof is standard and outlined for completeness in Appendix -D.
The sinusoidal portion of the sinogram $\text{SIN}_{x,y}$ should have enough information to recover the pixel intensity $(x,y)$ as it contains all the measurements associated with this pixel. Note however that the pixel at $(x,y)$ influences the integral over any line passing through it and thus also the parts of the sinogram corresponding to pixels on those lines. This can be loosely thought of as a consequence of non-orthogonality of the Radon transform. The above statement is thus more precisely a statement about the filtered sinogram since information is “relocalized” by the high-pass filtering step in the FBP.
This is related to the celebrated support theorems of Sigurdur Helgason, Jan Boman, and others [36, 37, 38, 39]. These theorems state that under appropriate conditions a compactly-supported image may be recovered from a compactly-supported subset of its Radon data. These results do not involve filtering explicitly, but its influence is implicit. They apply to idealized sampling and SNR conditions.
Indeed, the high-pass filtering in the FBP is derived for noiseless data and a continuum of observed angles. In reality the projections are corrupted with noise and come from a sparse subset of projection angles. We address this by 1) incorporating “contextual information” about the target pixel and 2) letting the filter be learnable to adapt it to the specifics of discretization and noise.
As shown in Figure 3, we exploit the spatial regularity of medical images (encoded in training data) by using the measurements which provide local information around $(x,y)$. This ensures that the model does not overfit large-scale features in the training data while maintaining low computational complexity. We thus additionally extract from the sinogram the regions associated with the neighboring pixels around $(x,y)$ and store this information in vector ${\mathbf{p}}_{x,y}$,
@@
where $K=C^{2}$ determines the number of neighboring pixels around $(x,y)$ for an odd number $C\geq 1$ and $d$ denotes the scale of the window which adjusts the receptive field. In order to recover the image at pixel $(x,y)$ from ${\mathbf{p}}_{x,y}$, we use a multi-layer perception $\text{MLP}_{\theta}:~{}\mathbb{R}^{r\times K}\to~{}\mathbb{R}$ parameterized by $\theta$,
$\hat{{\mathbf{f}}}(x,y)=\text{MLP}_{\mathrm{\bm{\theta}}}\big{[}{\mathbf{p}}_{x,y}\big{]},$ (9)
which estimates the pixel intensity $\hat{{\mathbf{f}}}_{x,y}$ from the local features around $(x,y)$. We call the proposed model Glimpse, standing for generalized [^1] The word “generalized” emphasizes that locality is also encoded in the transform domain, not just in real space as in some of earlier work. local imaging with MLPs. In the following section, we describe how our implementation of Glimpse allows to adapt to noisy measurements. We then propose a training strategy with resolution-agnostic memory usage in Section III-C. In Appendix -B, we show how Glimpse compensates for calibration errors. Further details for network architecture and training can be found in Appendix -A.
III-B Adaptive Filtering for Noisy Measurements

### v3

# Glimpse: Generalized Locality for Scalable and Robust CT

@@ Abstract
Deep learning has become the state-of-the-art approach to medical tomographic imaging. A common approach is to feed the result of a simple inversion, for example the backprojection, to a multiscale convolutional neural network (CNN) which computes the final reconstruction. Despite good results on in-distribution test data, this often results in overfitting certain large-scale structures and poor generalization on out-of-distribution (OOD) samples. Moreover, the memory and computational complexity of multiscale CNNs scale unfavorably with image resolution, making them impractical for application at realistic clinical resolutions. In this paper, we introduce Glimpse, a local coordinate-based neural network for computed tomography which reconstructs a pixel value by processing only the measurements associated with the neighborhood of the pixel. Glimpse significantly outperforms successful CNNs on OOD samples, while achieving comparable or better performance on in-distribution test data and maintaining a memory footprint almost independent of image resolution; 5GB memory suffices to train on 1024 $\times$ 1024 images which is orders of magnitude less than CNNs. Glimpse is fully differentiable and can be used plug-and-play in arbitrary deep learning architectures, enabling feats such as correcting miscalibrated projection orientations. Our implementation and Google Colab demo can be accessed at https://github.com/swing-research/Glimpse.
Keywords: Deep Learning, Computed Tomography, Image Reconstruction
1 Introduction
@@
Convolutional neural networks (CNNs) have become the standard approach for tomographic image reconstruction [1]. The U-Net [2] architecture underpins numerous deep learning reconstruction methods, achieving strong results on a variety of imaging problems including computed tomography (CT) [3], magnetic resonance imaging (MRI) [4] and photoacoustic tomography [5]. Its success is often attributed to the particular multi-scale architecture [6].
Despite remarkable progress with CNN-based methods, some core practical challenges complicate their application to real problems:
  • Poor Generalization under Distribution Shift: CNNs show good performance on in-distribution test images similar to the training data but tend to overfit class-specific image content. This results in poor robustness to distribution shift in data and sensing [7, 8]. Model-based networks address this drawback by integrating the forward and adjoint operators into multiple network layers or iterations [9, 10, 11, 12, 13, 14]. This, however, hurts scalability.
  • High Memory and Computation Cost: The required memory grows steeply with image resolution [15] for CNNs and even more steeply for model-based networks such as learned primal-dual (LPD) [10]. Moreover, unlike standard networks like U-Net which can handle large images by working on patches, model-based networks like LPD do not permit patch processing since the Radon transform in the network does not handle incomplete data.
Figure 1: Glimpse; $\text{NN}_{\theta}$ processes the measurements associated with the pixel $(x,y)$ and its neighbors extracted from the filtered sinogram. This local processing network has promising performance on OOD data while being computationally efficient all due to its locality.
1.1 Our Innovations
In this paper, we propose Glimpse, a novel coordinate-based local reconstruction framework for sparse-view CT. As shown in Figure 1, unlike large-scale CNNs that operate globally on filtered backprojection (FBP) [16] reconstructions, Glimpse estimates a given pixel value using only local measurements in the sinogram domain associated with this pixel. There is no backprojection step. Localization prevents Glimpse from overfitting the large-scale features and results in robust performance under distribution shift.
(a) Memory footprint (batch size 64)
(b) Training time (500 iterations)
Figure 2: The memory and time requirements during training vary across different models, with Glimpse being substantially faster and more memory-efficient compared to the baselines. Remarkably, Glimpse’s memory usage remains nearly constant regardless of image resolution, making it an excellent choice for high-dimensional image reconstruction tasks. All experiments were performed on a single A100 GPU with 80GB of memory. Missing data points indicate that the corresponding model exceeded the GPU’s memory capacity at the specified resolution.
At the same time, it results in high computational efficiency: the coordinate-based design permits training on mini-batches of both pixels and objects. This leads to fast and efficient training, requiring a small, fixed amount of memory almost independent from the image resolution. As shown in Figure 2, Glimpse requires significantly less memory and training time than CNNs, in particular compared with model-based networks like LPD. It can efficiently train on realistic images in resolution $1024\times 1024$ and beyond.
Glimpse is fully differentiable, all the way down to the sensing and integration geometry. This is an advantage over the standard CNN-based architectures. Most approaches to CT rely on fixed sensor geometry which is encoded in the forward operator, whether explicitly, as seen in methods like FBP [16], SART [17], LGS [9], and LPD [10] or implicitly in U-Net [2] when taking FBP as input. This fixed geometry is a problem when faced with uncertainties in calibration or blind inversion problems where the sensor geometry information is entirely unavailable [18, 19]. Our differentiable architecture allows us to estimate projection angles which results in better reconstructions. Furthermore, differentiability enables us to replace the fixed FBP filter by one that is optimal for the noise level and data distribution; this is illustrated in Figure 1. All this ultimately results in high-quality reconstructions.
1.2 Why are U-Nets Sensitive to Distribution Shift?
@@
Figure 4: Performance of U-Net [2] trained on chest images in resolution $128\times 128$: evaluation on in-distribution test data (chest samples) and OOD brain samples shows that the large receptive field of U-Net hinders its ability to generalize on OOD samples, with its PSNR even falling below that of FBP reconstruction. We indicate PSNRs between the reconstructions and the ground truth.
We close the introduction by presenting an experiment which illustrates why U-Net-like CNNs—which post-process FBP reconstructions—generalize poorly out-of-distribution. Figure 3 shows a point-like object, its sparse view sinogram, and the FBP reconstruction. It is evident that the FBP is supported over the entire field of view. This raises the question of the ideal receptive field size for CNNs like U-Net: a large receptive field is statistically beneficial to gather information correlated with the value of a target pixel [20, 21]. A similar argument shows that backprojection introduces long-range correlations in noise.
But the issue with models with large receptive fields is that they often overfit class-specific image content in training data which leads to poor generalization on out-of-distribution samples [22]. Indeed, Figure 4 shows that while U-Net produces good results when tested on chest images similar to training data, it performs poorly on out-of-distribution brain images. This is problematic in domains such as medical imaging where robustness over distribution shifts and other uncertain and variable factors is important [23].
2 Related Work
2.1 Model-based vs Model-free Inversion
@@
There are two major classes of deep learning to CT reconstruction: model-based and model-free. In the model-based approach, neural networks process raw sinograms and map them to the desired CT images while the Radon transform is integrated into multiple network layers or iterations [9, 10, 13, 24]. These methods perform remarkably well across various inverse problems, but they are computationally expensive, especially during training [15]. The high computational cost is due, among other factors, to the repeated application of the Radon transform and its adjoint in the network architecture.
@@
By contrast, model-free approaches offer a computationally cheaper alternative. The Radon transform (or its adjoint) is only used once in FBP computation before the neural network [3, 25, 26]. However, these models often require deep networks with a large receptive field to leverage the information delocalized across the FBP image. Recent studies aim to bypass the fixed FBP operator to provide greater flexibility. The common approach is a direct sinogram-to-image mapping that combines CNNs and MLP blocks, effectively replacing the FBP operator with learnable components [27, 28]. He et al. [29] present a partially learnable FBP by substituting the traditional Ram-Lak filter with an MLP block and incorporating learnable weighted averaging in the backprojection step. This modified FBP is further refined by a post-processing CNN. Recently, Hamoud et al. [21] used a measurement rearrangement technique to stratify backprojected features by angle and thus enable the use of smaller, shallower CNNs.
2.2 Robustness of deep learning for image reconstruction
As discussed in Section 1, deep neural networks often suffer from poor generalization and unstable reconstructions [8, 30, 7]. In [31], the authors present a theoretical study that highlights a trade-off between stability and accuracy and propose neural networks that navigate this trade-off and improve generalization. Genzel et al. study the role of network architecture in improving generalization [32]. Incorporating the forward operator and enforcing measurement consistency have been shown to substantially improve generalization [20, 10, 33]. Another technique to improve generalization is jittering by additive Gaussian noise during training [32, 34]. In this paper, we show that computationally efficient neural networks which incorporate the right notion of transform-domain locality achieve excellent generalization in- and out-of-distribution.
2.3 Implicit Neural Representation for Imaging
Glimpse is a coordinate-based reconstruction framework that recovers the image intensity at each pixel separately. Recently, neural fields, also known as implicit neural representations (INRs) [35, 36, 37], have emerged as a promising coordinate-based approach for representing continuous signals, images, and 3D volumes. Unlike traditional deep learning models that represent signals as discrete arrays, INRs use deep neural networks, typically MLPs, to map coordinates to signal values, enabling a continuous signal representation. This approach offers several advantages over conventional models. For instance, INRs can seamlessly interpolate signals within a continuous space instead of being limited to a single resolution. Moreover, their coordinate-based representation allows for flexible memory usage, making them particularly well-suited for high-dimensional 3D reconstructions [38, 39, 40, 41, 42, 43] and scene representations [44].
Coordinate-based models have also demonstrated strong performance in computational imaging. INRs efficiently model signals and their spatial derivatives which is useful for solving partial differential equations (PDEs) [35, 45]. They can be combined with self-supervised learning to learn a continuous representation of sub-sampled CT sinograms [46]. Zha et al. [47] use INRs to learn a continuous image representation that aligns with sinogram measurements for cone-beam CT reconstruction. Unlike all these methods, Glimpse learns a map from both measurements and coordinates to reconstruction values at individual pixels and is thus a true, learned image reconstruction operator rather than a signal parameterization.
2.4 Uncalibrated CT Imaging.
@@ 2.3 Implicit Neural Representation for Imaging
In CT imaging, the acquisition operator is usually known but only a limited number of measurements is collected, either to minimize radiation exposure or shorten acquisition time (sparse view) or when sample geometry and stage mechanics limit projection angles to a cone (limited view). In certain situations, the acquisition operator is only partially or approximately known. Neglecting this uncertainty can result in a significant drop in the quality of the reconstructions [18]. To tackle this challenge, total least squares approaches have been developed, involving the perturbation of an assumed forward operator [48, 49, 50] or trained networks combined with autodifferentiation and resampling [19].
3 Methods
In this section we introduce Glimpse. We begin with a brief overview of tomographic imaging in order to introduce the filtered backprojection formula.
3.1 Computed Tomography
@@
Tomographic imaging [51] plays an important role in many applications including medical diagnosis [52], industrial testing [53], and security [54]. We consider 2D computed tomography where the image of interest $f({\mathbf{x}})$ with size $D\times D$ is reconstructed from measurements of (X-ray) attenuation. The forward model is the Radon transform $Rf$ which computes integrals of $f({\mathbf{x}})$ along lines $L$,
@@
${\mathbf{s}}={\mathbf{A}}{\mathbf{f}}+{\mathbf{n}}$ (5)
@@
${\mathbf{f}}^{\text{FBP}}_{x,y}=\sum_{m=1}^{r}\tilde{{\mathbf{s}}}(y\cos(\alpha_{m})-x\sin(\alpha_{m}),m),$ (6)
where ${\mathbf{f}}^{\text{FBP}}\in\mathbb{R}^{N\times N}$ is the FBP reconstruction, $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$, ${\mathbf{h}}$ is a certain high-pass filter, * denotes the convolution and linear interpolation is used in (6) for evaluating $\tilde{{\mathbf{s}}}(x,\cdot)$ when $x$ is not an integer. As shown in Proposition 2 in Appendix .4, while the Ram-Lak filter is the optimal choice for ${\mathbf{h}}$ in the case of noise-free complete measurements, it amplifies noise in real measurements, yielding poor reconstructions.
With noise and an incomplete collection of projections, tomographic image reconstruction is an ill-posed inverse problem that requires an image prior as regularizer. We introduce our proposed method, Glimpse, designed to respect the geometry of CT, which implicitly learns such a prior from training data.
3.2 Glimpse: Generalized Local Imaging with MLPs
@@
$\text{SIN}_{x,y}(m)={\mathbf{s}}(y\cos(\alpha_{m})-x\sin(\alpha_{m}),m).$ (7)
Similar to (6), we can use interpolation to evaluate ${\mathbf{s}}(x,\cdot)$ for non-integer $x$. This localization is formally captured by the following proposition.
@@
The standard proof is outlined in Appendix .5.
This may seem to suggest that the neighborhood of the sinusoid-shaped part of the sinogram $\text{SIN}_{x,y}$ contains sufficient information to recover the pixel intensity at location $(x,y)$. Note however that the pixel at $(x,y)$ influences the integral over any line passing through it and thus also the parts of the sinogram corresponding to pixels on those other lines; this can be loosely thought of as a consequence of non-orthogonality of the Radon transform. The above statement is thus more accurately a statement about the filtered sinogram since high-pass filtering in the FBP “relocalizes” information. We mention in the passing that it is also related to the celebrated support theorems of Sigurdur Helgason, Jan Boman, and others [55, 56, 57, 58] which state that a compactly-supported image may be recovered from a compactly-supported subset of its Radon data under idealized sampling and SNR conditions.
Indeed, the high-pass filtering in the FBP is derived for noiseless data and a continuum of observed angles. In reality, the projections are corrupted with noise and come from a sparse subset of projection angles. We address this by 1) incorporating “contextual information” about the target pixel and 2) letting the filter be learnable to adapt it to the specifics of discretization and noise.
As shown in Figure 1, we exploit the spatial regularity of medical images (encoded in training data) by using the measurements that provide local information around $(x,y)$. This ensures that the model does not overfit large-scale features in the training data while maintaining low computational complexity. We thus additionally extract from the filtered sinogram the regions associated with the neighboring pixels around $(x,y)$ and store this information in vector ${\mathbf{p}}_{x,y}$,
@@
where $K=C^{2}$ determines the number of neighboring pixels around $(x,y)$ for an odd number $C\geq 1$ and $d$ denotes the scale of the window which adjusts the receptive field. In order to recover the image at location $(x,y)$ from ${\mathbf{p}}_{x,y}$, we use a neural network $\text{NN}_{\theta}:~{}\mathbb{R}^{r\times K}\to~{}\mathbb{R}$ with parameters $\theta$,
$\hat{{\mathbf{f}}}(x,y)=\text{NN}_{\mathrm{\bm{\theta}}}\big{(}{\mathbf{p}}_{x,y}\big{)},$ (9)
which estimates the pixel intensity $\hat{{\mathbf{f}}}_{x,y}$ from the local features around $(x,y)$. As we typically use a small neighborhood size $K$, we can parameterize $\text{NN}_{\theta}$ using a multi-layer perceptron (MLP). We call the proposed model Glimpse, standing for generalized [^1] The word “generalized” emphasizes that locality is also encoded in the transform domain, not just in real space as in some of earlier work. local imaging with MLPs. Glimpse can be viewed as a learnable alternative to FBP as it replaces the simple averaging along the corresponding sinusoidal support with a learnable non-linear operator, parameterized by $\text{NN}_{\theta}$, which processes the local contextual measurements. Our method can be seen as an interpolation between CNNs applied globally to FBP reconstructions and model-based architectures which explicitly employ the backprojection operator. This is because our inversion is structured ”like an FBP” (which simply sums filtered sinogram values along the sinusoidal support) whereas we allow for a more general function of the neighborhood of the sinusoidal support (and thus can approach optimal reconstruction for a larger class of priors than Gaussian processes).
In the following section, we provide further details regarding Glimpse’s architecture. We describe in Section 3.4 how our implementation of Glimpse allows adapting to noisy measurements. We then propose a training strategy with resolution-agnostic memory usage in Section 3.5. In Appendix .3, we show how backpropagating through Glimpse can compensate for calibration errors.
3.3 MultiMLP: efficient processing of increased projections
Figure 5: MultiMLP architecture; the input patch (here over a circular geometry) is split into smaller chunks each processed with a separate MLP, the extracted information is then mixed by another MLP. Each red point contains the associated sinusoidal curve extracted from the sinogram.
The number of parameters in $\text{NN}_{\theta}$ when parameterized by an MLP scales with the number of projections $r$ and the neighborhood size $K$, which increases computational complexity and slows down training. To mitigate this issue, we propose MultiMLP, a new architecture designed to efficiently process large numbers of projections and neighborhoods. Inspired by vision transformers [59], we partition the extracted measurements ${\mathbf{p}}_{x,y}$ into smaller chunks, each processed by a separate MLP, as illustrated in Figure 5. The outputs of these MLPs are then mixed using another MLP. For ease of visualization, we show a circular neighborhood where each red point represents its associated sinusoidal curve.
3.4 Adaptive Filtering for Noisy Measurements